# supplement_jasa

### v2

@@ 4.2 Simulation - Multivariate Transfer learning
Figure 6: Posterior credible intervals for parameters recovered by bps, with true parameters in yellow.
Table 2 presents predictive performance in terms of rmspe, along with the hyperparameter values corresponding to the different experimental settings. The first three rows correspond to fixed values of $\{\alpha,\phi\}$ in the Bayesian transfer learning approach of Section 2.1. The last row corresponds to bps when $\alpha$ and $\phi$ are weighted averages using the stacking weights as described in the last paragraph of Section 4.1. While bps outperforms transfer learning in terms of rmspe, some remarks on transfer learning are merited. The transfer learning approach offers slightly better predictive performance than the well-specified model with misspecified data than with well-specified data. This, while somewhat unexpected, is attributed to the estimated response surfaces being practically indistinguishable. Furthermore, transfer learning in even the highly misspecified setting performs competitively. In conclusion, these findings seem to be consistent with theoretical insights offered by Stein, (1988) and Stein and Handcock, (1989), who theoretically established the ability of Gaussian processes to deliver good predictive performance even for misspecified covariance functions within the in-fill asymptotic paradigm.
Model $\alpha$ $\phi$ Average rmspe
well specified 0.8 4 0.942
misspecified 0.45 6.63 0.963
highly misspecified 0.25 50 1.236
double bps 0.75 4.11 0.900
Table 2: Average rmspe and specification for each model scenario.
5 Application
Global warming and its critical consequences (see Fisher,, 1958; Nicholls,, 1989; Friehe et al.,, 1991; O’Carroll et al.,, 2019) are increasingly being investigated using machine learning and AI tools that assist in developing predictive global models and formulating data-driven policies. trying to produce faster solutions. The key aspects of global warming are related to global temperature, atmospheric and land surface compositions. We analyze data extracted from NASA’s Moderate Resolution Imaging Spectroradiometer (modis) platform.
Specifically, we consider as response variables two vegetation indices present in vi data: the normalized difference vegetation index (ndvi); and the red reflectance (rr). Both measure the activity of vegetation on land surfaces and are commonly investigated to help understand biophysical and structural properties of vegetation at global scales (Walther et al.,, 2018; Jiang et al.,, 2020; Ai et al.,, 2020). The predictors comprise an intercept and the solar zenith angle for that location ($p=2$). The latter depicts the strength of solar irradiation. All the data were collected over a 16-day window in May 2024 and then averaged.
[…]
@@ 5 Application
The multivariate model developed in Section 2.3 with ndvi and rr comprising the $q=2$ columns of $Y$. Figures 7 and 8 illustrate maps corresponding to ndvi and rr, respectively, using $K=2,000$ subsets. The top left panel in each figure presents the spatially interpolated map of the training data for the respective responses. The interpolated maps for the test data are displayed in the top right panel of the respective figures. These maps reveal pronounced spatial variation, where the darker shades of green in ndvi represent higher values of detected biomass, while lighter shades of brown represent low biomass. Conversely for rr, warmer colors in the red spectrum represent higher reflectance.
@@
Figure 8: Left to right: Maps for training data (top left), test data (top right) and predicted surface (bottom right) for rr. Empirical coverage for held-out values of outcomes (bottom left). Results correspond to $K=2,000$.
Figures 7 and 8 show results from double bps. Maps of interpolated posterior means and predicted responses are displayed in the bottom right of the respective figures. The bps interpolated maps are almost indistinguishable from the true surface (top right). This suggests that the automated bps effectively, perhaps even strikingly, recovers the spatial patterns in spite of the modeling simplifications over more elaborate statistical models (Banerjee,, 2020; Zhang et al.,, 2021). Moreover, the double bps behaves well in terms of predictive performance and uncertainty quantification for both values of $K$ considered. It achieves an rmspe of $(0.206,0.159)$ when $K=2,000$, and $(0.241,0.197)$ with $K=4,000$ subsets for ndvi and rr respectively. This is significantly lower (by almost 100%) than the standard Bayesian conjugate multivariate linear model, which produces rmspe of $(0.474,0.439)$ for the two vegetation indexes, respectively. The predictive $95\%$ credible intervals when $K=2,000$ for both responses are shown in the bottom left of Figures 7 and 8, respectively.
[…]
Conjugate Linear model bps ($K=4,000$) bps ($K=2,000$)
@@
$\beta_{1,\textsc{ndvi}}$2.708 ( -2.719, -2.697) 0.744 ( 0.478, 0.998) 0.707 ( 0.485, 0.939)
$\beta_{0,\textsc{rr}}$16.924 (-17.014, -16.837) -0.418 (-2.470, 1.710) -0.562 (-2.187, 0.962)
@@
$\Sigma_{\textsc{ndvi},\textsc{rr}}$0.167 ( -0.168, -0.166) -0.114 (-0.173, -0.079) -0.076 (-0.094, -0.061)
@@
$\varrho_{\textsc{ndvi},\textsc{rr}}$0.921 ( -0.928, -0.915) -0.816 (-0.874, -0.706) -0.844 (-0.880, -0.795)
rmspe [0.474, 0.404, 0.439] [0.241, 0.197, 0.219] [0.206, 0.159, 0.183]
Time 12 21 81
Table 3: Vegetation Index data analysis parameter estimates, rmspe, and computing time in minutes for candidate models. Parameter posterior summary 50 (2.5, 97.5) percentiles. Root mean square prediction error(s) presentation [ndvi, red reflectance, average].
Recalling that all the analyses are produced on a personal laptop (with just 5 physical cores) with minimal human intervention, the total run time of only 21 minutes with $K=4,000$ and 81 minutes with $K=2,000$ are impressive confirm the quadratic dependence of the partition size discussed in Section LABEL:sec:data1. Moreover, the strong dependence on the number of $J$ competitive models assessed is worth noting. Here $J=4$ while in Section LABEL:sec:data1 $J=3$. This suggests a marginal computational burden for each competitive model.
@@ 6 Discussion
This manuscript devises a statistical modeling component for an artificially intelligent geospatial system (GeoAI) tailored to analyze massive datasets with minimal human intervention. The contribution harnesses analytically accessible multivariate statistical distributions in conjunction with Bayesian predictive stacking to deliver rapid inference on multivariate processes and circumventing iterative algorithms that require, often extensive, human tuning or control. Our proposed GeoSpatial AI system is capable of handling truly massive amounts of spatial data that preclude accessibility even on high performance CPU architectures and we develop Bayesian transfer learning using double Bayesian predictive stacking.
@@
A few additional remarks are warranted. The development here has been elucidated with a hierarchical matrix-variate spatial process framework. While modeling simplifications have been introduced to minimize human intervention, we wish to emphasize that the bps seamlessly applies to more versatile but analytically intractable models. First, for spatially misaligned multivariate data where not all the variables may have been observed over the same set of locations, we encounter missing entries in $Y$ but are still able to deliver analytically tractable closed-form distribution theory (Zhang and Banerjee,, 2022). Second, should users wish to explore more complex models with unknown $\{\alpha,\phi\}$ or work with alternate multivariate models, then we can implement full Bayesian inference (now using MCMC, Variational Bayes or Laplace approximations) on each subset and invoke bps to pool across the subsets of the data. Time for analyzing each subset increase due to model complexity, but scaling inference to massive datasets is still enabled by bps—and that too with considerable efficiency over alternate divide-and-conquer approaches such as the geometric median, or gm, of posteriors (used in meta-kriging developed by Guhaniyogi and Banerjee,, 2018).
@@
A preliminary examination of Section 2.3 reveals that posterior inference is conceptualized as a “mixture of mixtures”. Following Yao et al., (2018), we prefer stacking to fitting a mixture model because the former is numerically robust and requires almost no human tuning. Moreover, the structure of mixture-of-mixtures conceptually resembles the “Mixture of Experts” (MoE) that are adopted by AI platforms such as GPT-4 and Mistral. Disseminating our proposed product with our accompanying software, which is currently being migrated to R, is expected to significantly boost Geospatial AI systems. Future directions will also explore perceived potential of bps as a feeder for emerging amortized inference methods (Ganguly et al.,, 2023; Zammit-Mangion et al.,, 2024; Sainsbury-Dale et al.,, 2024) to achieve Bayesian inference. Rapid delivery of posterior estimates of the entire spatial process from bps will amount to more training data for amortized neural learners that can result in accelerated tuning for subsequent Bayesian inference. We do not see our proposed approach as a competitor to, but rather as supplementary to amortized neural inference. Such developments will be pursued as future research. We also seek to expand and fully investigate automated double bps using Markovian graphical structures across data subsets to further expedite and improve Geospatial AI systems.
@@
The online supplement includes additional derivations and details on matrix-variate distributions (Section LABEL:sec:distributiontheory), computational details and algorithms (Section LABEL:sec:compdetails), additional simulation experiment results (Section LABEL:sec:sim_suppl), details on accelerated bps for univariate spatial data (Section LABEL:sec:AccelerUnivarMod) and analysis of an additional dataset on sea surface temperature (Section LABEL:sec:data1). Section LABEL:sec:eda outlines automated computation of model hyperparameters without human intervention. Computer programs to reproduce all experiments and data analysis are publicly accessible from the GitHub repository lucapresicce/Bayesian-Transfer-Learning-for-GeoAI

### v3

@@ 4.2 Amortized Bayesian inference
Figure 4: Amortized posterior credible intervals for parameters. True parameters in yellow.
Figure 5: Surface interpolations for true spatial process, BPS prediction (50 quantile), and Amortized prediction of $\{50,2.5,97.5\}$ quantiles. Each row corresponds to an outcome.
5 Data analysis
5.1 Vegetation index data
Statisticians and scientists face growing demands to analyze and study global warming datasets (see, e.g., Fisher, , 1958; Nicholls, , 1989; Friehe et al., , 1991; O’Carroll et al., , 2019). The sheer volume of such datasets and the process-driven models required for their analysis have naturally raised the question of migrating such analysis to AI platforms. Our current application analyzes vegetation index data from the Moderate Resolution Imaging Spectroradiometer (modis) developed by nasa. We specifically focus on “mod13c1.061 - Terra Vegetation Indices 16-Day L3 Global 0.05 Deg Climate Modeling Grid” (Didan, , 2021), which provides vegetation indices per pixel on a 0.05-degree climate modeling grid (3600 rows by 7200 columns of 5600-meter pixels). It contains 16-day global composites, cloud-contamination-free, with additional reflectance and angular information.
Modeling Normalized Difference Vegetation Index (ndvi) and Red Reflection (rr) jointly is scientifically important because it separates and defines the specific biophysical factors influencing vegetation. ndvi is built on the fundamental principle that healthy vegetation absorbs red light and reflects near-infrared (NIR) light. By examining both the calculated NDVI and its red reflectance component, scientists can gain deeper insights into vegetation’s structural and biochemical properties. The solar zenith angle (sza), which is the angle between the Sun and the point directly overhead, is a shared predictor that influences how much solar irradiation reaches the surface and is, therefore, crucial when assessing biomass and vegetation indices. ndvi and rr reflect the ability of vegetation to absorb photosynthetically active radiation and help scientists understand processes underlying climate change (Tucker, , 1979; Sellers, , 1985; Justice et al., , 1985; Haque et al., , 2024).
Vegetation index Mean Std.Dev Min. Max. Histogram Boxplot
ndvi 8.593 0.517 6.909 9.469 [Uncaptioned image] [Uncaptioned image]
Red Reflectance 8.563 0.447 8.007 9.472 [Uncaptioned image] [Uncaptioned image]
Table 1: Summary statistics and visual representation of response variables.
The dataset comprises $1,500,000$ locations. We randomly selected $1,002,500$ sites, reserving $2,500$ for predictive evaluation. Following Zhang and Banerjee, (2022), outcomes were log-transformed and we labeled $\log(\textsc{ndvi}+1)$ as ndvi and $\log(\textsc{rr}+1)$ as rr. All variables, including sza, were averaged over a 16-day window in May 2024. Table 1 summarizes the spatial responses distribution. The maximum inter-site distance among sites is approximately $42,909$ kilometers.
Our central scientific objective is to jointly predict ndvi and rr from massive, globally distributed datasets of this scale using GeoAI. We assess whether multivariate statistical models can deliver accurate, scalable, and timely predictions across millions of spatial locations while accounting for dependence structures in the solar zenith angle. These capabilities are crucial for application in ecology, agriculture, and climate policy.
5.2 Results using dbps
To analyze our Vegetation Index (vi) data, we follow machine-generated exploratory data analysis (eda) described in Section LABEL:sec:eda. The predictors comprise an intercept and the solar zenith angle for that location ($p=2$). Based on Section LABEL:sec:eda, we set $\alpha\in\{0.825,0.909\}$, and $\phi\in\{0.049,0.067\}$ respectively. We specify $\{\gamma,\Sigma\}$ in (2.5) using $m_{0}=0_{p\times q}$, $M_{0}=10\mathbb{I}_{p}$, $\Psi_{0}=\mathbb{I}_{q}$, $\nu_{0}=3$, and finally we opt for an exponential spatial covariance function. Following insights from Section LABEL:sec:sim5_sens, we fix the subset size at $n_{k}\in\{250,500\}$, leading to a number of subsets $K\in\{4,000\;,\;2,000\}$ respectively. We use a random scheme to form the partition of the dataset and present results when $K=2,000$ and in Table 2 for $K=4,000$.
@@ 5.2 Results using dbps
The multivariate model developed in Section 2.3 with ndvi and rr comprising the $q=2$ columns of $Y$. Figures 6 and 7 illustrate maps corresponding to ndvi and rr, respectively, using $K=2,000$ subsets. The top left panel in each figure presents the spatially interpolated map of the training data for the respective responses. The interpolated maps for the test data are displayed in the top right panel of the respective figures. These maps reveal pronounced spatial variation, where the darker shades of green in ndvi represent higher values of detected biomass, while lighter shades of brown represent low biomass. Conversely for rr, warmer colors in the red spectrum represent higher reflectance.
@@
Figure 7: Left to right: Maps for training data (top left), test data (top right) and predicted surface (bottom right) for $r$. Empirical coverage for held-out values of outcomes (bottom left). Results correspond to $K=2,000$.
Figures 6 and 7 show results from double bps. Maps of interpolated posterior means and predicted responses are displayed in the bottom right of the respective figures. The dbps interpolated maps are almost indistinguishable from the true surface (top right). This suggests that the automated dbps effectively, perhaps even strikingly, recovers the spatial patterns in spite of the modeling simplifications over more elaborate statistical models (Banerjee, , 2020; Zhang et al., , 2021). The predictive $95\%$ credible intervals when $K=2,000$ for both responses are shown in the bottom left of Figures 6 and 7, respectively.
Table 2 compares model parameters posterior estimates for dbps with $K\in\{4,000\;,\;2,000\}$, and the Bayesian conjugate linear model, which does not take into account spatial variability. The notably higher magnitudes of the intercepts in the non-spatial linear model are unsurprising as the spatial random effects absorb much of the variation in the intercepts. The solar zenith angle is positively associated with rr and negatively associated with ndvi in the non-spatial model. This, too, is expected since higher levels of solar irradiation are associated with higher red reflectance and with more arid regions with less vegetation. However, the spatial models reveal that the slope for ndvi is significantly positive with solar irradiation after the spatial effects have absorbed previously unaccounted latent or lurking factors that might have contributed to the negative slopes in the non-spatial model.
Parameter Conjugate Linear model dbps (K=4, [^0] K=4{,}000) dbps (K=2, [^0] K=2{,}000)
@@
$\beta_{1,\textsc{ndvi}}$ -2.708 (-2.719, -2.697) 0.744 (0.478, 0.998) 0.707 (0.485, 0.939)
$\beta_{0,\textsc{rr}}$ -16.924 (-17.014, -16.837) -0.418 (-2.470, 1.710) -0.562 (-2.187, 0.962)
@@
$\Sigma_{\textsc{ndvi},\textsc{rr}}$ -0.167 (-0.168, -0.166) -0.114 (-0.173, -0.079) -0.076 (-0.094, -0.061)
@@
Table 2: Vegetation Index data analysis parameter estimates for candidate models. Parameter posterior summary 50 (2.5, 97.5) percentiles.
Model Time (min) RMSPE $\varrho_{\textsc{ndvi},\textsc{rr}}$
dbps (K=2, [^0] K=2{,}000) 81 [0.206, 0.159, 0.183] -0.844 (-0.880, -0.795)
dbps (K=4, [^0] K=4{,}000) 21 [0.241, 0.197, 0.219] -0.816 (-0.874, -0.706)
Conjugate Linear Model 12 [0.474, 0.404, 0.439] -0.921 (-0.928, -0.915)
gbm 2 [0.419, 0.352, 0.385] -0.993 (–,–)
drf 10 [0.420, 0.353, 0.386] -0.991 (–,–)
dnn 26 [0.422, 0.354, 0.388] -0.987 (–,–)
automl 44 [0.419, 0.352, 0.385] -0.993 (–,–)
Table 3: Vegetation Index data analysis computing time in minutes, rmspe, and empirical correlation ($\varrho$) for candidate models. Root mean square prediction error(s) presentation [ndvi, red reflectance, average].
In addition to double bps (dbps) and Bayesian multivariate linear regression, we expanded the analysis by including four competitive algorithms: distributed random forest (drf), gradient boosting (gbm), deep neural network (dnn), and a fully automatic machine learning algorithm (automl). As detailed in Section LABEL:sec:compdetails, we implement and perform the analysis using the h2o R package (Fryda et al., , 2024). As no method for multivariate models is available within the h2o framework, we fit these algorithms separately for ndvi, and rr, considering the same explanatory variables. Then, we compute the empirical correlation among the predictions. In Table 3, we offer a comparison of computational costs, predictive performances, and empirical correlation. The double bps dominates both running time and predictive performances. It achieves an rmspe of $(0.206,0.159)$ when $K=2,000$, and $(0.241,0.197)$ with $K=4,000$ subsets for ndvi and rr respectively. This is significantly lower (by almost 100%) than the standard Bayesian conjugate multivariate linear model, which produces rmspe of $(0.474,0.439)$ for the two vegetation indexes, respectively. Actually, the double bps outperformed all the other competitive algorithms in terms of rmspe. This is attained at a fraction of the computational time for other methods, let alone double bps models multivariate distributions jointly, and offer full explainability. Estimates of $\varrho_{\textsc{ndvi}}$ and $\varrho_{\textsc{rr}}$ reveal a well-documented negative association between the two responses. Indeed, the spatial patterns in these indices are almost the reverse of each other as revealed in Figures 6 and 7. The conjugate Bayesian linear model estimates higher negative correlation, while dbps tends to underestimate it, considerably so as $n_{k}$ decreases. We attempted comparisons with the lmc and bart models, but omitted due they exceeded memory.
Recalling that all the analyses are produced on a personal laptop (with just 5 physical cores) with minimal human intervention, the total run time of only 21 minutes with $K=4,000$ and 81 minutes with $K=2,000$ for double bps are impressive and confirm the quadratic dependence of the partition size discussed in Section LABEL:sec:sim5_sens. Moreover, the strong dependence on the number of $J$ competitive models assessed is worth noting. Here $J=4$, this suggests a marginal computational burden for each competitive model.
@@ 6 Discussion
This manuscript devises a statistical modeling component for an artificially intelligent geospatial system (GeoAI) tailored to analyze massive data with minimal human intervention. The contribution harnesses analytically accessible multivariate statistical distributions in conjunction with Bayesian predictive stacking to deliver rapid inference by circumventing iterative algorithms that require, often extensive, human tuning. Our proposed GeoSpatial AI system relies on Bayesian transfer learning using double Bayesian predictive stacking to process massive amounts of streaming spatial data on high-performance CPU architectures.
@@
Some additional remarks are warranted. The development here has been elucidated with a hierarchical matrix-variate spatial process framework. Although modeling simplifications have been introduced to minimize human intervention, we emphasize that dbps seamlessly applies to more versatile but analytically intractable models. For spatially misaligned multivariate data, where not all variables have been observed in the same set of locations, we encounter missing entries in $Y$ using analytically tractable closed-form distribution theory (Zhang and Banerjee, , 2022). Should users wish to explore more complex models with unknown $\{\alpha,\phi\}$ or work with alternate multivariate models, we can implement Bayesian inference (using MCMC, Variational Bayes or Laplace approximations) on each subset and invoke bps to pool across the subsets of the data. The time for analyzing each subset increases due to model complexity, but scaling inference to massive datasets is still achieved by bps—and this too with considerable efficiency over divide-and-conquer approaches such as the geometric median, or gm, of posteriors (see meta-kriging in Guhaniyogi and Banerjee, , 2018). Future research can build upon recent work by Cabel et al., (2025) to enrich spatial-temporal dependence structures within dbps and to further accelerate amortized inference.
@@
Section 2.3 devised posterior inference as a “mixture of mixtures”. Following Yao et al., (2018), we prefer stacking to fitting a mixture model because the former is numerically robust and requires almost no human tuning. Moreover, the structure of mixture-of-mixtures conceptually resembles the “Mixture of Experts” (MoE) that are adopted by AI platforms such as GPT-4 and Mistral. Disseminating our proposed product with our accompanying software, which is currently being migrated to R, is expected to significantly boost Geospatial AI systems. Future directions will also explore perceived potential of dbps as a feeder for emerging amortized inference methods (Ganguly et al., , 2023; Zammit-Mangion et al., , 2024; Sainsbury-Dale et al., , 2024) to achieve Bayesian inference. Rapid delivery of posterior estimates of the entire spatial process from dbps will amount to more training data for amortized neural learners that can result in accelerated tuning for subsequent Bayesian inference. We do not see our proposed approach as a competitor to, but rather as supplementary to amortized neural inference. Refer to Section 4.2 for an effective representation. Such developments will be pursued as future research. We also seek to expand and fully investigate automated double bps using Markovian graphical structures across data subsets to further expedite and improve Geospatial AI systems.
@@
The online supplement includes exploratory data analyses, and automated settings for model hyperparameters without human intervention (Section LABEL:sec:eda), additional derivations and details on matrix-variate distributions (Section LABEL:sec:distributiontheory), some theoretical insights (Section LABEL:sec:upperbound), computational details and algorithms (Section LABEL:sec:compdetails), and additional simulation experiment results (Section LABEL:sec:sim_suppl_M) Computer programs to reproduce all our analysis are publicly accessible from the GitHub repository lucapresicce/Bayesian-Transfer-Learning-for-GeoAI